# Vision-based Manipulation from Single Human Video with Open-World Object Graphs

### v1

@@ 3.3 Robot Policy To Synthesize Actions
Given a manipulation plan, ORION constructs a policy that synthesizes actions, detailed in Figure 3. The manipulation policy is derived based on the aforementioned learning objective to achieve object motion similarities. The action synthesis comprises three major steps: identify a keyframe from the plan that matches the current observation, predict object motions, and use the predictions to optimize the robot actions for the robot controller to execute. The policy repeats these three steps until a task is completed or fails, detailed in Appendix C.
@@
Retrieving OOGs from the plan. ORION identifies the keyframe and retrieves OOGs to help decide what next actions to take. At test-time, ORION localizes objects in the new observations and estimates contact relations using the same vision pipeline as described in Section 3.2. Then ORION retrieves the OOG that has the same set of relations as the current state, allowing us to identify a pair $(\mathcal{G}_{l},\mathcal{G}_{l+1})$, where $\mathcal{G}_{l}$ is the retrieved graph and $\mathcal{G}_{l+1}$ the graph of the next keyframe. This pair of graphs provides sufficient information to decide which object to manipulate next, termed the target object, and we denote its point cloud at keyframe $l$ as $\hat{o}_{\text{target}}$, and its keypoint trajectories as $\tau^{V}_{\text{target}}$. A target object is the one in motion due to manipulation between two keyframes, and it is determined by computing the average velocity per-object using motion features in $G_{l}$. At the same time, another object, called the reference object, is involved in changing contact state relations from $\mathcal{G}_{l}$ to $\mathcal{G}_{l+1}$ and serves as a spatial reference for the target object’s movement. We use the point cloud of the reference object at next keyframe $l+1$, as the reference object might have location changes due to object interactions and using the updated information from the next keyframe gives us an accurate prediction of the trajectories. Once the target and reference objects are determined, we can localize the corresponding objects in the new observations and their point clouds are denoted as $o_{\text{target}}$ and $o_{\text{ref}}$, respectively.
@@
Predicting object motions. Given the target and reference objects from keyframes $l$, and $l+1$, we predict the motion of the target object in the current state by warping the keypoint trajectories estimated from $V$. To warp the trajectories, we first identify the initial and goal locations of keypoints in the new configuration by leveraging information given by the OOG pair. We use global registration of point clouds [25] to align $\hat{o}_{\text{target}}$ with $o_{\text{target}}$ and $\hat{o}_{\text{ref}}$ with $o_{\text{ref}}$, giving us two transformations to compute the new starting and goal positions of target object keypoints conditioned on where the reference object is. Then we normalize $\tau^{V}_{\text{target}}$ with its starting and goal locations, obtaining $\hat{\tau}_{\text{target}}$. $\hat{\tau}_{\text{target}}$ only contains the directional and curvature patterns that are independent of the absolute location of the initial and the goal keypoints. Then we scale it back to the workspace coordinate frame using the new starting and goal locations, resulting in new keypoint trajectories of the target object $\tau^{Ro}_{\text{target}}$.
@@
$\min_{T_{0},T_{1},\dots,T_{t_{l+1}-t_{l}}}\sum_{i=0}^{t_{l+1}-t_{l}}(\tau^{Ro}_{\text{target}}({i+1})-T_{i}\tau^{Ro}_{\text{target}}({i}))$ (1)
where $\tau^{Ro}_{\text{target}}(i)$ $(0\leq i\leq t_{l+1}-t_{l})$ represents the keypoint locations at timestep $i$ along the trajectory. This optimization process naturally allows generalizations over spatial variations, as the action sequence always conditions on a new location instead of overfitting to absolute locations. To further specify where the gripper should interact with the object and whether it should be open or closed, we augment the resulting SE(3) sequence with the interaction information stored in the hand node $h$. We implement a combination of inverse kinematics (IK) and joint impedance control to achieve precise and compliant execution.
@@ 4 Experiments
In this section, we report on experiments to answer the following questions regarding the effectiveness of ORION and the important design choices. 1) Is ORION effective at constructing manipulation policies given a single human video in the open-world setting? 2) To what extent does the object-centric abstraction improve the policy performance? 3) How critical is it to model the object motions with keypoints and the TAP formulation? 4) How consistent is the performance of ORION’s policy given videos taken in different conditions? 5) How effectively does ORION scale to long-horizon manipulation tasks?
@@ 4.1 Experiment Setup
Task descriptions. We design the following five tasks to evaluate the policy performance: 1) Mug-on-coaster: placing a mug on the coaster; 2) Simple-boat-assembly: putting a small red block on a toy boat; 3) Chips-on-plate: placing a bag of chips on the plate; 4) Succulents-in-llama-vase: inserting succulents into the llama vase; 5) Rearrange-mug-box: placing a mug on a coaster and placing a cream cheese box on a plate consecutively; 6) Complex-boat-assembly: placing both a small red block and a chimney-like part on top of a boat. 7) Prepare-breakfast: placing a mug on a coaster and putting a food box and can on the plate. The first four are “short-horizon” tasks, and the last three are “long-horizon” tasks. In the context of this paper, “short-horizon” refers to tasks that only require one contact relation between two objects, while “long-horizon” refers to those that require more than one contact relation. Detailed success conditions of all tasks are described in Appendix C.
@@
Experimental setup. We design experiments to fully test the efficacy of our method by providing the robot with videos captured in everyday scenarios, which naturally encompass visual backgrounds and camera setups that are different from the one for the robot. Specifically, we record an RGB-D video of a person performing each of the five tasks in everyday scenarios, such as an office or a kitchen. We use an iPad for recording, which comes with a TrueDepth Camera, and we fix it on a camera stand. The videos can be found in the supplementary materials. During test time, the robot receives visual data through a single RGB-D camera, Intel Realsense435, and performs manipulation in its workstation to evaluate policies. We use the 7DoF Franka Emika Panda robot for all the experiments.
@@
Evaluation protocol. As we describe in the experimental setup, the videos naturally include various visual backgrounds and camera perspectives that are significantly different from the robot workspace. Therefore, we only intentionally vary two dimensions before evaluating each trial of robot execution, namely the spatial layouts and the new object instances. Furthermore, the new object generalizations are included in the tasks Mug-on-coaster and Chips-on-plate as mugs and chip bags have many similar instances. As for the other three tasks, there are no novel objects involved, but we extensively vary the spatial layouts of task-relevant objects for evaluation. The policy performance of a task is the averaged success rates over 15 real-world trials. Aside from the success rates, we also group the failed executions into three types: Missed tracking of objects due to failure of the vision models, Missed grasping of objects during execution, and Unsatisifed contacts where the target object configurations are not achieved for reasons other than the previous two failure types.
@@
Baselines. To understand the model capacity and validate our design choices, we compare ORION with baselines. Since no prior work exists that matches the exact setting of our approach, we adopt the most important components from prior works and treat them as baselines to our model. Specifically, we implement the following two baselines: 1) Hand-motion-imitation [9, 26] is a baseline that predicts robot actions by learning from the hand trajectories. The rest of the parts remain the same as ORION. We use this baseline to show whether it is critical to compute actions centering around objects. 2) Dense-Correspondence [15, 27] is a baseline that replace the TAP model in ORION with a dense correspondence model, optical flows. This baseline is used to evaluate whether our choice of TAP model is a better design. For this ablative study, we conduct experiments on Mug-on-coaster and Simple-boat-assembly to validate our model design, covering the distribution of common daily objects and assembly manipulation that requires precise control.
@@ 4.2 Experimental Results
Our evaluations are presented in Figures 4111Individuals in images have been digitally blurred to ensure anonymity in accordance with the requirements of the double-blind review process. and  5. We answer question (1) by showing the successful deployment of the ORION policies, while no other methods are designed to be able to operate in our setting. Furthermore, ORION yields an average of $69.3\%$ success rates, which validates our model design in learning from a single human video in the open-world setting.
@@
To answer question (4), we conduct controlled experiments using the task Mug-on-coaster. Specifically, we record two additional videos of the same task in very different visual conditions and spatial layouts (see details in Appendix D) and construct a policy from each video. Then, we compare the two policies against the original one and test them using the same set of evaluation conditions. The results in Figure 5(b) shows that there is no statistically significant difference in the performance, demonstrating that ORION is robust to videos taken under very different visual conditions. Finally, we show that ORION is effective in scaling to long-horizon tasks. This conclusion is supported by the performance among the pairs of Mug-on-coaster versus Rearrange-mug-box, and Simple-boat-assembly versus Complex-boat-assembly. In these two pairs, both the short-horizon tasks are subgoals of their long-horizon counterparts, yet we do not see any performance drop between the two. This result indicates that ORION excels at scaling to long-horizon tasks without a significant drop in policy performance.
@@ 5 Related Work
Learning Manipulation From Human Videos. Human videos offer a rich repertoire of object interaction behaviors, making them an invaluable data source for manipulation. A large body of work has explored how to leverage human video data for learning robot manipulation [9, 10, 28, 29, 30, 31, 32, 33], either through pre-training a single latent representation [7, 9, 33], learning an implicit reward function [6, 8], or learning generative models that in-paint human morphologies [15, 26, 28, 34]. However, they either require additional robot data from the target tasks or paired data between humans and robots. Our approach takes a novel direction by tackling how a robot can imitate or learn from a single human video only: the robot does not rely on pre-existing data, models, or ground-truth annotations in scenes where video recording and robot evaluation take place. We refer to such a setting as open-world imitation from observation, where the robot is not programmed or trained to interact with the objects in the video a priori and the video data does not come with any robot actions. Our setting is closely related to the problem of “Imitation Learning from Observation” [12], where state-only demonstrations are used to construct policies for physical interaction. However, this line of prior work assumes simulators of demonstrated tasks exist and physical states of the agents or objects are known [35, 36, 37, 38, 39]. In contrast, our setting does not assume the digital replica of real-world tasks, and all the object information is only perceived through RGB-D videos.
@@
Learning Manipulation From a Single Demonstration. Studies have delved into learning manipulation policies from one demonstration. A notable frame is one-shot imitation learning within meta-learning framework proposed by Duan et al. [40]. While prior works on one-shot imitation learning have shown a robot performing new tasks from one demonstration, they require extensive in-domain data and a well-curated set of meta-training tasks beforehand, leading to significant data collection costs and restricted policy generalization at test time due to the tailored nature of the training.
@@
An alternative approach involves using a single demonstration for initial guidance, refining the policy through real-world self-play [41, 42, 43, 44, 45]. However, this approach mainly applies to reset-free tasks and struggles with scaling to multi-stage tasks where resetting to the task initial conditions does not come free. Our work aligns with these studies in using a single demonstration for learning manipulation, but stands out by not needing prior data or self-play. With just one single human video, our method constructs a policy that successfully completes the task, adapting to various visual and spatial differences from the task instance of video demonstration.
@@
Object-Centric Representation for Learning Robot Manipulation. The concept of object-centric representation has long been recognized for its potential to enhance robotic perception and manipulation by focusing on the objects within a scene. Prior works have shown effectiveness of such representation in downstream manipulation tasks by factorizing visual scenes into disentangled object concepts [46, 47, 48, 49, 50], but these works are typically confined to known object categories or instances. Recent developments in foundation models allow robots to access the open-world object concepts through pre-trained vision models [13, 14], enabling a wide range of abilities such as imitation of long-horizon tabletop manipulation [5, 51] or mobile manipulation in the wild [52]. Building upon these advances, our work focuses on leveraging open-world, object-centric concepts in imitating manipulation behaviors from actionless human videos. We propose a graph-based representation called Open-world Object Graph (OOG), which allows a robot to imitate from a human video by leveraging the object-centric concepts. This proposed representation shares a similar vein with prior works that factorize scene or task-relevant visual concepts into scene graphs [29, 53, 54, 55, 56]. However, our representation is tailored to integrate open-world object concepts and enable generalization across different embodiments, specifically a human and a robot.

### v2

@@ 3.3 Robot Policy To Synthesize Actions
Given a manipulation plan, ORION constructs a policy that synthesizes actions, detailed in Figure 3 and Algorithm  2. The manipulation policy is derived based on the aforementioned learning objective to achieve object motion similarities. The action synthesis comprises three major steps: identify a keyframe from the plan that matches the current observation, predict object motions, and use the predictions to optimize the robot actions for the robot controller to execute. The policy repeats these three steps until a task is completed or fails. Task completion conditions are detailed in Appendix C.
@@
Retrieving OOGs from the plan. ORION first identifies the keyframe and retrieves OOGs to help decide what next actions to take (Lines 2-4). At test-time, ORION localizes objects in the new observations and estimates contact relations using the same vision pipeline as described in Section 3.2. Then, ORION retrieves the OOG that has the same set of relations as the current state, allowing us to identify a pair $(\mathcal{G}_{l},\mathcal{G}_{l+1})$, where $\mathcal{G}_{l}$ is the retrieved graph and $\mathcal{G}_{l+1}$ the graph of the next keyframe. This pair of graphs provides sufficient information to decide which object to manipulate next, termed the target object, and we denote its point cloud/pose at keyframe $l$ as $\hat{o}_{\text{target}}$, and its object-centric feature trajectories as $\tau^{V}_{\text{target}}$ (Lines 5-6). A target object is the one in motion due to manipulation between two keyframes. In the RGB-D case, the target object is identified by computing the average velocity per-object using motion features in $G_{l}$. At the same time, another object, called the reference object, is involved in changing contact state relations from $\mathcal{G}_{l}$ to $\mathcal{G}_{l+1}$ and serves as a spatial reference for the target object’s movement. In the RGB case, ORION uses the outputs of the classification done by the VLM to determine which objects are target and reference objects. In both cases, ORION use the point cloud of the reference object at the next keyframe $l+1$, as the reference object might have location changes due to object interactions, and using the updated information from the next keyframe gives us an accurate prediction of the trajectories. Once the target and reference objects are determined, we can localize the corresponding objects in the test-time observations, where their point clouds/poses are denoted as $o_{\text{target}}$ and $o_{\text{ref}}$, respectively.
@@
Predicting object motions. Given the target and reference objects from keyframes $l$ and $l+1$, ORION predicts the motion of the target object in the current state by warping the object-centric feature trajectories estimated from $V$ (Lines 7-8). To warp the trajectories, ORION first identifies the initial and goal locations of keypoints in the new configuration by leveraging information given by the OOG pair. In the RGB-D case, ORION uses global registration of point clouds [28] to align $\hat{o}_{\text{target}}$ with $o_{\text{target}}$ and $\hat{o}_{\text{ref}}$ with $o_{\text{ref}}$, giving two transformations to compute the new starting and goal positions of target object keypoints conditioned on where the reference object is. In the RGB case, since we have the 6-DOF poses of $\hat{o}_{\text{target}}$, $o_{\text{target}}$, $\hat{o}_{\text{ref}}$, and $o_{\text{ref}}$, ORION directly finds the relative transformations required to align the respective pairs.
Now, given new starting and goal positions of the target objects, ORION must warp the original object-centric feature trajectory to fit these new start and end conditions. In the RGB-D case, ORION is warping a 3D point trajectory, which we call translation warping (full details in Algorithm 3). ORION first normalizes $\tau^{V}_{\text{target}}$ with its starting and goal locations, obtaining $\hat{\tau}_{\text{target}}$. $\hat{\tau}_{\text{target}}$ only contains the directional and curvature patterns that are independent of the absolute location of the initial and the goal keypoints. Then, $\hat{\tau}_{\text{target}}$ is scaled back to the workspace coordinate frame using the new starting and goal locations, resulting in new keypoint trajectories of the target object $\tau^{Ro}_{\text{target}}$.
In the RGB case, since we have a 6-DOF pose trajectory, ORION also needs to do rotation warping in addition to translation warping (full details in Algorithm 4). In rotation warping, ORION  first computes the alignment rotation $A$ that maps the demonstration’s reference‑to‑target object direction to the test-time reference-to-target object direction, and left-multiplies $A$ on every rotation in $\tau^{V}_{\text{target}}$. ORION then normalizes the resulting sequence about its first frame—so it begins at the identity rotation—yielding a “relative rotation” trajectory.
Next, ORION scales this normalized sequence back to the workspace by defining two boundary rotations: one that maps the normalized start to the real initial pose, and one that maps the normalized end to the real goal pose. ORION uses SLERP between these two rotations to get a smooth transition $\{I_{i}\}$, and post‑multiplies the normalized rotations by $I_{i}$, at each timestep. This set of operations guarantees the warped rotation begins and ends exactly at the real‑scene boundary conditions, while still containing the same rotation trajectory “shape” from the original $\tau^{V}_{\text{target}}$.
@@
$\operatorname*{arg\,min}_{T_{0},T_{1},\dots,T_{t_{l+1}-t_{l}}}\sum_{i=0}^{t_{l+1}-t_{l}}\left\lVert\tau^{Ro}_{\text{target}}({i+1})-T_{i}\tau^{Ro}_{\text{target}}({i})\right\rVert_{2}^{2}$ (1)
where $\tau^{Ro}_{\text{target}}(i)$ $(0\leq i\leq t_{l+1}-t_{l})$ represents the keypoint locations at timestep $i$ along the trajectory (Line 9) and $t_{l}$ represents the time index for the keyframe $l$. This optimization process naturally allows generalizations over spatial variations, as the action sequence always conditions on a new location instead of overfitting to absolute locations. In the RGB case, however, because ORION already has the warped 6-DOF pose trajectory of the object, we no longer need to solve Equation 1 and can find $T_{i}$ in the 6-DOF object pose trajectory.
To further specify where the gripper should interact with the object and whether it should be open or closed, ORION augments the resulting SE(3) sequence with the interaction information stored in the grip node $h$ (Line 11). We implement a combination of inverse kinematics (IK) and joint impedance control to achieve precise and compliant execution (Lines 10, 12).
@@ 4 Experiments
In this section, we report on experiments to answer the following questions regarding the effectiveness of ORION and the important design choices:
  1. 1.
    Is ORION effective in constructing manipulation policies from single human videos in the open-world setting?
  2. 2.
    To what extent do the object-centric abstractions improve the policy performance?
  3. 3.
    How critical is it to model the object motions with keypoints and the TAP formulation?
  4. 4.
    How consistent is ORION’s performance on videos taken in various conditions?
  5. 5.
    Can ORION scale to multi-step, long-horizon manipulation tasks?
  6. 6.
    Can ORION effectively construct manipulation policies from RGB-only demonstration videos in the absence of depth information?
@@ 4.1 Experiment Setup
Task descriptions. We design the following seven tasks to evaluate the policy performance:
  1. 1.
    Mug-on-coaster: placing a mug on the coaster.
  2. 2.
    Simple-boat-assembly: putting a small red block on a toy boat.
  3. 3.
    Chips-on-plate: placing a bag of chips on the plate.
  4. 4.
    Succulents-in-llama-vase: inserting succulents into the llama vase.
  5. 5.
    Pour-juice: pouring juice into a container.
  6. 6.
    Peas-on-plate: putting a can of peas on a plate.
  7. 7.
    Cheese-on-plate: putting a block of cheese on a plate.
  8. 8.
    Rearrange-mug-box: placing a mug on a coaster and placing a cream cheese box on a plate consecutively.
  9. 9.
    Complex-boat-assembly: placing both a small red block and a chimney-like part on top of a boat.
  10. 10.
    Prepare-breakfast: placing a mug on a coaster and putting a food box and can on the plate.
@@
The first seven are “short-horizon” tasks, and the last three are “long-horizon” tasks. In the context of this paper, “short-horizon” refers to tasks that only require one contact relation between two objects, while “long-horizon” refers to those that require more than one contact relation. Finally, we note that Mug-on-coaster and Succulents-in-llama-vase have both RGB and RGB-D demonstration videos, while Pour-juice, Peas-on-plate, and Cheese-on-plate only have RGB demonstration videos.
Detailed success conditions of all tasks are described in Appendix C.
@@
Experimental setup. We design experiments to fully test the efficacy of our method by providing the robot with videos captured in everyday scenarios, which naturally encompass visual backgrounds and camera setups that are different from those of the robot. Specifically, we record an RGB-D video of a person performing each of the seven tasks in everyday scenarios, such as an office or a kitchen. We use an iPad for recording, which comes with a TrueDepth Camera, and we fix it on a camera stand. The videos can be found on the project website. During test time, the robot receives visual data through a single RGB-D camera, Intel RealSense D435, and performs manipulation in its workstation to evaluate policies. We use the 7-DoF Franka Emika Panda robot for all the experiments.
@@
Evaluation protocol. As we describe in the experimental setup, the videos naturally include various visual backgrounds and camera perspectives that are significantly different from the robot workspace. Therefore, we only intentionally vary two dimensions before evaluating each trial of robot execution, namely the spatial layouts and the new object instances. Furthermore, the new object generalizations are included in the tasks Mug-on-coaster, Chips-on-plate, Pour-juice, Peas-on-plate, and Cheese-on-plate. As for the other tasks, there are no novel objects involved, but we extensively vary the spatial layouts of task-relevant objects for evaluation. The policy performance of a task is the average success rate over 15 real-world trials. Aside from the success rates, we also group the failed executions into three types: Missed tracking of objects due to failure of the vision models, Missed grasping of objects during execution, and Unsatisfied contacts where the target object configurations are not achieved for reasons other than the previous two failure types.
Figure 5: This figure shows results in the same format as Figure 4, with the difference that demonstration videos do not contain depth information and that videos come from different sources. Note that the Mug-on-coaster task uses a different demonstration video to showcase compatibility with different object instances, camera angles, and video sources.
@@
Baselines. To understand the model capacity and validate our design choices, we compare ORION with baselines. Since no prior work exists  [^1] See Section 5 for detailed discussion on related work that matches the exact setting of our approach, we adopt the most important components from prior works and treat them as baselines to our model. Specifically, we implement the following two baselines: 1) Hand-motion-imitation [9, 29] is a baseline that predicts robot actions by learning from the hand trajectories. The rest of the parts remain the same as ORION. We use this baseline to show whether it is critical to compute actions centered around objects. 2) Dense-Correspondence [16, 30] is a baseline that replace the TAP model in ORION with a dense correspondence model, optical flows. This baseline is used to evaluate whether our choice of the TAP model is a better design. For this ablative study, we conduct experiments on Mug-on-coaster and Simple-boat-assembly to validate our model design, covering the distribution of common daily objects and assembly manipulation that requires precise control.
@@ 4.2 Experimental Results
Our evaluations are presented in Figures 4  5, and  6. We answer question (1) by showing the successful deployment of the ORION policies, while no other methods are designed to be able to operate in our setting. Furthermore, ORION yields an average of $66.7\%$ success rate, which validates our model design in learning from a single human video in the open-world setting.
@@
To answer question (4), we conduct controlled experiments using the task Mug-on-coaster. Specifically, we record two additional videos of the same task in very different visual conditions and spatial layouts (see details in Appendix D) and construct a policy from each video. Then, we compare the two policies against the original one and test them using the same set of evaluation conditions. The results in Figure 6(b) show that there is no statistically significant difference in the performance, demonstrating that ORION is robust to videos taken under very different visual conditions. Next, to answer question (5), we show that ORION is effective in scaling to long-horizon tasks. This conclusion is supported by the performance among the pairs of Mug-on-coaster versus Rearrange-mug-box, and Simple-boat-assembly versus Complex-boat-assembly. In these two pairs, both the short-horizon tasks are subgoals of their long-horizon counterparts, yet we do not see any performance drop between the two. This result indicates that ORION excels at scaling to long-horizon tasks without a significant drop in policy performance.
Finally, we answer question (6) by observing a similar successful deployment of ORION’s policies when demonstration videos only contain RGB information without depth, as shown in Figure 5. Pour-juice is retrieved from an in-the-wild YouTube video, Peas-on-plate and Mug-on-coaster are generated from Veo 2 [15], and Succulents-in-llama-vase and Cheese-on-plate are from the same recordings used for RGB-D experiments but with the depth channel removed. We observe an average success rate of $85.3\%$, showing that our method can perform well even without depth information. In addition, we see that our method is able to handle demonstration videos from different sources (from controlled environments, Internet videos, and generated models) and different camera angles, along with test-time object instances and camera angles that differ from the demonstration. These results support our conclusions for question (4) in the RGB-only setting. To directly compare the RGB-only and RGB-D pipelines, we inspect the results of the Succulents-in-llama-vase task, on which both pipelines were evaluated. Interestingly, we see that the RGB-only pipeline performs comparably and even slightly better than the RGB-D pipeline on the same task. We note that certain RGB-only architectural changes, such as the shift from 3D keypoint trajectories to 6-DOF pose trajectories, allow the model to trade noise in global registration and end-effector trajectory optimization for noise in object pose estimation and mesh generation models. We observe that the strong world knowledge embedded in these models allows our RGB-only pipeline to perform competitively with our RGB-D pipeline, compensating for the lack of depth with minimal added noise.
@@ 5 Related Work
Learning Manipulation From Human Videos. Human videos offer a rich repertoire of object interaction behaviors, making them an invaluable data source for manipulation. A large body of work has explored how to leverage human video data for learning robot manipulation [9, 10, 31, 32, 33, 34, 35, 36], either through pre-training a single latent representation [7, 9, 36], learning an implicit reward function [6, 8], or learning generative models that in-paint human morphologies [16, 29, 31, 37]. However, they either require additional robot data from the target tasks or paired data between humans and robots. Our approach takes a different direction by tackling how a robot can imitate or learn from a single human video only: the robot does not rely on pre-existing data, models, or ground-truth annotations in scenes where video recording and robot evaluation take place. We refer to such a setting as open-world imitation from observation, where the robot is not programmed or trained to interact with the objects in the video a priori and the video data does not come with any robot actions. Our setting is closely related to the problem of “Imitation Learning from Observation” [12], where state-only demonstrations are used to construct policies for physical interaction. However, this line of prior work assumes simulators of demonstrated tasks exist and physical states of the agents or objects are known [38, 39, 40, 41, 42]. In contrast, our setting does not assume the digital replica of real-world tasks, and all the object information is perceived only through videos.
@@
One-Shot Imitation Learning for Manipulation. Studies have delved into learning manipulation policies from one demonstration. A notable frame is one-shot imitation learning within a meta-learning framework proposed by Duan et al. [43]. While prior works on one-shot imitation learning have shown a robot performing new tasks from one demonstration, they require extensive in-domain data and a well-curated set of meta-training tasks beforehand, leading to significant data collection costs and restricted policy generalization at test time due to the tailored nature of the training.
@@
An alternative approach involves using a single demonstration for initial guidance, refining the policy through real-world self-play [44, 45, 46, 47, 48]. However, this approach mainly applies to reset-free tasks and struggles with scaling to multi-stage tasks where resetting to the task’s initial conditions does not come free. Our work aligns with these studies in using a single demonstration for learning manipulation, but stands out by not needing prior data or self-play. With just one single human video, our method constructs a policy that successfully completes the task, adapting to various visual and spatial differences from the task instance of video demonstration.
Open-World Imitation Learning in Robotics. Recent work [49, 50] has begun to bridge single-demonstration learning and learning from human videos, closely matching our “open-world” setting. Similarly to ORION, these works utilize the broad world knowledge of off-the-shelf models. However, they typically rely on stronger geometric inputs (e.g., calibrated RGB-D or multi-view captures, explicit 3D hand poses, or pre-scanned object meshes) that are rarely available in casual consumer or Internet videos. These dependencies limit applicability to uncontrolled, single-view recordings. In contrast, our method requires only the demonstration video (RGB, or optionally RGB-D) with no pre-scanned object models.
Another recent approach, OKAMI [51], most directly addresses the problem of “open-world imitation from observation”. However, OKAMI and ORION represent fundamentally different approaches to imitation. OKAMI’s method is demonstrator-centric, focusing on reconstructing the demonstrator’s body and hand motions and retargeting them onto a humanoid robot. In contrast, our method is object-centric, focusing on generating a plan that replicates the observed object trajectories, independent of the human’s specific pose. In addition, this means that ORION does not need a demonstration video showing the human’s entire body. By extracting object-centric Open-world Object Graphs (OOGs) with off-the-shelf vision foundation models, ORION recovers task-relevant object states and interaction cues directly from the video and constructs robot-executable plans that generalize across backgrounds, viewpoints, spatial layouts, and novel object instances.
@@
Object-Centric Representations for Robot Manipulation. The concept of object-centric representation has long been recognized for its potential to enhance robotic perception and manipulation by focusing on the objects within a scene. Prior works have shown the effectiveness of such representation in downstream manipulation tasks by factorizing visual scenes into disentangled object concepts [52, 53, 54, 55, 56], but these works are typically confined to known object categories or instances. Recent developments in foundation models allow robots to access the open-world object concepts through pre-trained vision models [13, 14], enabling a wide range of abilities such as imitation of long-horizon tabletop manipulation [5, 57] or mobile manipulation in the wild [58]. Building upon these advances, our work focuses on leveraging open-world, object-centric concepts in imitating manipulation behaviors from actionless human videos. We propose a graph-based representation called Open-world Object Graph (OOG), which allows a robot to imitate from a human video by leveraging the object-centric concepts. This proposed representation shares a similar vein with prior works that factorize scene or task-relevant visual concepts into scene graphs [32, 59, 60, 61, 62]. However, our representation is tailored to integrate open-world object concepts and enable generalization across different embodiments, specifically a human and a robot.